### v3

# Navigate the Unknown: Enhancing LLM Reasoning with Intrinsic Motivation Guided Exploration

Abstract
Reinforcement learning (RL) has emerged as a pivotal method for improving the reasoning capabilities of Large Language Models (LLMs). However, prevalent RL approaches such as Proximal Policy Optimization (PPO) and Group-Regularized Policy Optimization (GRPO) face critical limitations due to their reliance on sparse outcome-based rewards and inadequate mechanisms for incentivizing exploration. These limitations result in inefficient guidance for multi-step reasoning processes. Specifically, sparse reward signals fail to deliver effective or sufficient feedback, particularly for challenging problems. Furthermore, such reward structures induce systematic biases that prioritize exploitation of familiar trajectories over novel solution discovery. These shortcomings critically hinder performance in complex reasoning tasks, which inherently demand iterative refinement across ipntermediate steps. To address these challenges, we propose an Intrinsic Motivation guidEd exploratioN meThOd foR LLM Reasoning (i-MENTOR), a novel method designed to both deliver dense rewards and amplify explorations in the RL-based training paradigm. i-MENTOR introduces three key innovations: trajectory-aware exploration rewards that mitigate bias in token-level strategies while maintaining computational efficiency; dynamic reward scaling to stabilize exploration and exploitation in large action spaces; and advantage-preserving reward implementation that maintains advantage distribution integrity while incorporating exploratory guidance. Experiments across three public datasets demonstrate i-MENTOR’s effectiveness with a 22.39% improvement on the difficult dataset Countdown-4. The source code is available for reference  [^1] https://github.com/Gaojingtong/iMENTOR.
1 Introduction
Reinforcement learning (RL) [2, 32, 17] has become essential for training Large Language Models (LLMs) [15, 24], evolving from Proximal Policy Optimization (PPO) [31] to Group-Regularized Policy Optimization (GRPO) [33]. Frameworks like DeepSeek [14, 20, 33] showcase State-Of-The-Art (SOTA) implementations that optimize multi-step reasoning using trajectory-level advantage estimation. RL uniquely transforms sparse rewards [12] into learning gradients, enabling coherent Chain-of-Thought (CoT) [40, 28] reasoning. By refining action distributions through feedback, RL bridges the gap between linguistic competence and goal-directed problem-solving in LLMs.
While contemporary RL methods like PPO and GRPO demonstrate progress through training with outcome-based rewards [36, 19, 37], two fundamental limitations persist for reasoning tasks. First, the reliance on static reward functions creates a sparse learning signal [3, 6] that fails to guide intermediate reasoning steps, forcing models to navigate vast action spaces with terminal outcome feedback alone. This is particularly evident on difficult samples and datasets [22] where outcome-based rewards are more likely to be zero. Second, despite GRPO’s group-wise sampling strategy, the inherent reward structure disincentivizes genuine exploration - trajectories yielding identical final outcomes receive equivalent advantage estimates regardless of different reasoning CoTs, effectively penalizing computationally intensive sampling efforts [46, 21]. This creates a paradoxical scenario where models optimize for reward exploitation at the expense of systematic exploration, particularly detrimental in difficult reasoning tasks that require complicated reasoning processes.
Traditional exploration methods (e.g., RND[5], ICM[27], Count-Based Exploration [25]) encourage agents to explore novel or under-visited states via intrinsic rewards, inspired by cognitive theories of curiosity. While effective in hard exploration environments, these methods, though promising for guiding LLM exploration, face challenges in LLM reasoning tasks due to: 1) Dynamic episodic length and computational overload arise due to dynamic CoT length and token-level exploration rewards [5]. Directly using each reasoning step as a sample for traditional exploration methods will result in the exploration reward of the long sequence being higher than that of the short sequence, inducing the model to explore samples with longer sequences. Moreover, the per-token computation for lengthy LLM outputs incurs significant computational overhead. 2) Large action space poses a challenge due to the exponential growth in possible reasoning paths for LLMs [48, 41]. Naive exploration strategies become computationally prohibitive, and the majority of randomly sampled trajectories fail to produce meaningful outcomes. 3) The conflict of the exploration rewards and the outcome rewards arises when exploration rewards are directly incorporated, disrupting outcome-based advantage and value estimation in methods such as PPO [31] and GRPO [33]. This introduces noise, destabilizing policy learning and diminishing the effectiveness of reward normalization.
In this paper, we propose an Intrinsic Motivation guidEd exploratioN meThOd foR LLM Reasoning (i-MENTOR), which mitigates reward sparsity, enables smart explorations in LLM training, and effectively resolves the above challenges: 1) Trajectory-aware exploration rewards operate at the sequence level to eliminate computational overload and sequence-length bias caused by dynamic episodic length. By employing two lightweight trajectory-aware networks, this component efficiently captures reasoning sequence uniqueness while maintaining computational traceability. 2) Dynamic reward scaling integrates: (i) Exploration reward regularization to stabilize training by mitigating reward scale fluctuations, (ii) error-conditioned reward trigger that selectively activates exploration rewards exclusively on incorrect reasoning trajectories, ensuring efficient explorations especially on hard samples, and (iii) a policy-preserving exploration attenuation on exploration rewards that adaptively reduces exploration intensity during the training. These techniques enable efficient exploration in the large action space while maintaining a stable, adaptive exploration-exploitation trade-off throughout policy optimization. 3) Advantage-preserving reward implementation introduces exploration incentives after advantage computation. This approach resolves inherent conflicts between exploration rewards and outcome-based rewards. By preserving the statistical integrity of outcome-driven advantage distributions, it integrates exploratory guidance to enable effective coordination between these objectives.
By unifying above mechanisms, i-MENTOR generates dense, stable, and computationally efficient exploration rewards that seamlessly integrate into RL-based LLM training pipelines. This enables enhanced reasoning capabilities through balanced exploration-exploitation dynamics, overcoming the limitations of conventional exploration rewards in LLM-based sequential decision-making contexts. Our contributions are summarized as follows:
  • We propose i-MENTOR, a novel method that systematically incorporates dense, stable, and computationally efficient exploration rewards to enhance LLM reasoning. Designed for seamless integration with established RL algorithms, i-MENTOR leverages exploration rewards to improve learning efficiency while maintaining stable policy updates throughout the training process.
  • We present three advances: (1) Trajectory-aware rewards generate exploration rewards while maintaining efficiency; (2) Dynamic reward scaling stabilizes exploration-exploitation with reward scaling approaches; (3) Advantage-preserving implementation injects exploration rewards after advantage computation to preserve policy gradient fidelity. Together, i-MENTOR systematically mitigates reward sparsity and enhances LLM reasoning through coordinated explorations.
  • Experiments on three public datasets demonstrate i-MENTOR’s consistent effectiveness across both PPO and GRPO, significantly enhancing the reasoning capabilities of LLMs.
2 Method
In this section, we first briefly introduce the current SOTA RL training methods for LLM, i.e., PPO and GRPO, then introduce our method, i-MENTOR.
2.1 Preliminary
2.1.1 Proximal Policy Optimization (PPO)
Proximal Policy Optimization (PPO) enhances policy optimization through a clipped objective function that ensures stable updates:
$\mathcal{J}_{\mathrm{PPO}}(\theta)=\mathbb{E}_{(q,a)\sim\mathcal{D},o_{\leq t}\sim\pi_{\theta_{\text{old }}}(\cdot\mid q)}\left[\min\left(w_{t}(\theta)\hat{A}_{t},\operatorname{clip}\left(w_{t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{t}\right)\right]$ (1)
where $w_{t}(\theta)=\frac{\pi_{\theta}\left(o_{t}\mid q,o_{<t}\right)}{\pi_{\theta_{\text{old }}}\left(o_{t}\mid q,o_{<t}\right)}$. Here, $(q,a)$ denotes question-answer pairs from dataset $\mathcal{D}$. $o_{t}$ and $o_{<t}$ are generated responses end at token position $t$ and $t-1$. The current and previous policies are parameterized by $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ respectively. The advantage estimator $\hat{A}_{t}$ employs Generalized Advantage Estimation (GAE) [30] with outcome-based reward function $R$ and value function $V$ for advantage estimation, while $\varepsilon$ controls the clipping range. By treating question-response sequences ending at each token position of $o_{t}$ as distinct states and constraining policy updates within a trust region, PPO stabilizes RL training for LLMs. However, the joint optimization of policy and value functions introduces computational overhead, limiting training efficiency.
2.1.2 Group-Regularized Policy Optimization (GRPO)
Group-Regularized Policy Optimization (GRPO) extends PPO through group-wise advantage normalization, formulated as:
$\mathcal{J}_{\mathrm{GRPO}}(\theta)$ $=\mathbb{E}_{(q,a)\sim\mathcal{D},\left\{o_{i}\right\}_{i=1}^{G}\sim\pi_{\theta_{\text{old }}(\cdot\mid q)}}$ (2)
${\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\left|o_{i}\right|}\sum_{t=1}^{\left|o_{i}\right|}\left(\min\left(w_{i,t}(\theta)\hat{A}_{i},\operatorname{clip}\left(w_{i,t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{i}\right)-\beta D_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\right)\right].}$
where $w_{i,t}(\theta)=\frac{\pi_{\theta}\left(o_{i,t}\mid q,o_{i,<t}\right)}{\pi_{\theta_{\text{old }}}\left(o_{i,t}\mid q,o_{i,<t}\right)}$, $\hat{A}_{i}=\frac{R_{i}-\operatorname{mean}\left(\left\{R_{i}\right\}_{i=1}^{G}\right)}{\operatorname{std}\left(\left\{R_{i}\right\}_{i=1}^{G}\right)}$, $G$ responses $\{o_{i}\}_{i=1}^{G}$ are generated per input, $\pi_{\mathrm{ref}}$ denotes the reference policy, and $\beta$ controls KL penalty strength. GRPO replaces value function estimation with group-wise normalization on outcome-based rewards, improving training efficiency and effects through reduced computational complexity and group-wise sampling process.
Figure 1: i-MENTOR. Values with green/yellow boxes denote higher outcome-based rewards/ advantages and exploration rewards; red/blue denote lower values. Black boxes mark the excluded values. $\hat{A}_{old}$ and $\hat{A}_{new}$ denote advantages derived from outcome-based rewards and i-MENTOR.
2.2 i-MENTOR
Despite advancements, PPO and GRPO face two key limitations: (1) Sparse reward signals that provide limited training guidance, and (2) weak exploration mechanisms during policy updates. These issues hinder learning efficiency and final optimization. To address these challenges, we propose an Intrinsic Motivation guidEd exploratioN meThOd foR LLM Reasoning (i-MENTOR), which introduces structured exploration to enhance LLM reasoning and can be easily implemented to SOTA RL methods like PPO and GRPO. As shown in Figure 1, vanilla RL methods with outcome-based rewards focus only on correct trajectories, ignoring under-visited ones and causing LLMs to converge suboptimally. i-MENTOR instead rewards incorrect under-visited trajectories, preventing the model from getting stuck via intrinsic motivation guided exploration.
A Case Study is presented in Figure 2 to illustrate the significance of i-MENTOR. Specifically, vanilla GRPO-trained LLM exhibits logical errors (e.g., missing numbers, redundancy in ①), repetitive reasoning patterns (②) due to intensive exploitations without explorations, and failures to solve tasks (③) or apply critical operations like multiplication (④). With the guidance of the exploration rewards from i-MENTOR, LLM trained with i-MENTOR-GRPO reduces logical errors, diversifies reasoning paths (④), and results in a correct answer (③), demonstrating improved reasoning ability. Additionally, as shown in Figure 3(a), i-MENTOR increases the average response length during training, stabilizing at a higher level than vanilla RL methods. This suggests that i-MENTOR ’s exploration strategies enable LLMs to learn more complex reasoning processes with longer CoTs through considering diverse reasoning paths in responses, leading to performance gains. In the following subsections, we will systematically introduce i-MENTOR through detailing its key components.
Figure 2: Case study on Countdown-4 [11] with simplified responses. Numbers with circles (e.g., ①) highlight key differences. Detailed case study with complete responses is provided in Appendix A.
(a) Response length.
(b) RND input statistics.
(c) Exploration rewards.
Figure 3: Training statistics on Countdown-4 [11].
2.2.1 Trajectory-aware Exploration Rewards
We propose i-MENTOR to provide intrinsic exploration rewards for LLM reasoning based on the Random Network Distillation (RND) framework [5] as it is computationally efficient and does not need prior knowledge. The original RND formulation employs a fixed randomly initialized target network $f_{\theta_{T}}$ to generate fixed outputs, together with a trainable predictor network $f_{\theta_{P}}$ (with identical architecture) that minimizes prediction error on visited states:
$\mathcal{L}(\boldsymbol{s}_{t})=\|f_{\theta_{P}}(\boldsymbol{s}_{t})-f_{\theta_{T}}(\boldsymbol{s}_{t})\|_{2}^{2}$ (3)
where $\boldsymbol{s}_{t}$ denotes the state at time step $t$ in reinforcement learning. The predictor network is updated at each batch or step in coordination with the main RL algorithm, such as PPO. This setup allows the prediction error to act as an intrinsic reward, encouraging the agent to explore novel states that exhibit high model uncertainty (i.e., large $\mathcal{L}(\boldsymbol{s}_{t})$). The exploration reward is thus defined as:
$R^{rnd}(\boldsymbol{s}_{t})=\frac{\|f_{\theta_{P}}(\boldsymbol{s}_{t})-f_{\theta_{T}}(\boldsymbol{s}_{t})\|_{2}^{2}}{\operatorname{std}\left(\|f_{\theta_{P}}(\boldsymbol{s}_{t})-f_{\theta_{T}}(\boldsymbol{s}_{t})\|_{2}^{2}\right)}.$ (4)
The denominator employs an exponential moving average of standard deviations during training updates to maintain reward magnitude consistency and temporal stability, while remaining fixed during evaluation for deterministic behavior.
However, for LLM reasoning where states correspond to tokens $\boldsymbol{s}_{t}=[q,o_{t}]$, direct RND application introduces two challenges: (1) Dynamic episodic length: Sequences $(q,o)$ with dynamic CoT length in response $o$ generate inputs $\{\boldsymbol{s}_{t}\}_{t=1}^{|o|}$ with lengths $|q|+1$ to $|q|+|o|$ for RND update. This results in systematic under-sampling of longer input sequences, as they only appear in extended responses, while shorter inputs recur across all sequences regardless of response length, as illustrated in Figure 3(b). (2) Computational overload: Processing thousands of tokens per sequence requires excessive RND network updates.
These phenomena are particularly evident in Figure 3(b). In this figure, a total of $253+416+515+549+701=2434$ RND inputs are generated from merely 5 rollout responses of 1 question with length $[253,416,515,549,701]$ in GRPO optimization. Shorter input states dominate these samples, inducing lower exploration rewards from frequent visits. RND requires multiple forward passes per sequence, creating computational overload.
To address these challenges, we propose trajectory-aware exploration rewards operating at the sequence level. Our approach maintains the dual-network architecture but processes complete sequences $(q,o)$ instead of multiple token-level states $\{\boldsymbol{s}_{t}\}_{t=1}^{|o|}$ as inputs for predictor / target networks:
[…]
Here $\mathcal{L}(q,o)$ is the loss function for updating $f_{\theta_{P}}$ while $R^{\star}_{1}(q,o)$ provides a single exploration reward per sequence. This design eliminates token-level bias through uniform sequence treatment and reduces computational complexity from $O(|o|)$ to $O(1)$ per training sequence, enabling stable and efficient LLM training.
2.2.2 Dynamic Reward Scaling
Despite the benefits of trajectory-aware exploration rewards, the large action space in LLM reasoning introduces challenges for training. Naive exploration strategies become infeasible due to reward range instability and suboptimal exploration efficiency. Specifically, we observe that indiscriminate application of exploration rewards leads to excessive fluctuations in the reward range, insignificant exploration effects, and performance bottlenecks in the later stage of model training. To address these challenges, i-MENTOR implements three dynamic scaling mechanisms:
  • Exploration reward regularization: Standard deviation-based scaling fails to contain reward magnitude fluctuations that destabilize training. In this paper, we instead employ min-max scaling within each batch for stable reward ranges:
    {$&r(q,o)=\|f_{\theta_{P}}(q,o)-f_{\theta_{T}}(q,o)\|_{2}^{2}\\ &R^{\star}_{2}(q,o)=\alpha\cdot\frac{r(q,o)-\min\limits_{(q,o)\in B}(r(q,o))}{\max\limits_{(q,o)\in B}(r(q,o))-\min\limits_{(q,o)\in B}(r(q,o))}\end{aligned}\right.$ (6)
    where $B$ denotes the training batch and hyperparameter $\alpha$ controls reward intensity. This regularization prevents the predictor network convergence from distorting the exploration reward distribution, maintaining stable learning signals throughout training.
  • Error-conditioned reward trigger: Uniform reward across correct and incorrect samples reduces exploration efficiency. To explore more effectively, we argue that more attention should be paid to samples that are not completely correct in order to both facilitate the discovery of diverse ideas and improve the overall performance efficiently. Therefore, i-MENTOR applies an error-conditioned reward trigger for different sampling sequences:
    $R^{\star}_{3}(q,o)=I_{a\neq o}\cdot R^{\star}_{2}(q,o)$ (7)
    where $a$ represents the ground truth answer. This conditioning directs exploration resources toward challenging samples with higher error potential, improving reward utilization efficiency.
  • Policy-preserving exploration attenuation: Unrestricted exploration hinders policy refinement in later training stages. We introduce step-based decay to balance exploration-exploitation tradeoffs:
    $R^{\star}(q,o)=\frac{\gamma}{\gamma+n}\cdot R^{\star}_{3}(q,o)$ (8)
    where hyperparameter $\gamma$ modulates decay rate with training step $n$. This schedule prioritizes early-stage exploration while gradually shifting focus to exploitation, enabling stable policy convergence.
As shown in Figure 3(c), during initial training steps, the predictor network rapidly converges, causing i-MENTOR to stop assigning high exploration rewards to non-novel responses, which sharply reduces the average reward. Three mechanisms further optimize the exploration rewards in the subsequent training steps: (1) the application of exploration reward regularization ($R^{\star}_{2}(q,o)$) prevents further decay despite the predictor network convergence; (2) the application of error-conditioned reward trigger deprecates rewards for correct ones, prioritizing efficient exploration on difficult samples ($R^{\star}_{2}(q,o)$); and (3) the application of policy-preserving exploration attenuation ($R^{\star}(q,o)$) gradually shifts focus from exploration to exploitation over time. These dynamic reward scaling approaches help balance exploration and exploitation, improving the overall training effect.
2.2.3 Advantage-preserving Reward Implementation
Since the intensity of the exploration behavior varies at different stages of training, directly combining exploration rewards with outcome-based rewards creates conflicting learning signals across training stages. For PPO, this injects noise into value function estimation. For GRPO, exploration rewards may invert originally positive advantage signs after group normalization, which is a critical issue since exploration rewards should never penalize samples. To resolve this conflict and ensure seamless integration with established RL algorithms like PPO and GRPO, i-MENTOR applies exploration rewards after advantage computation and possible value estimation through:
$\hat{A}_{new}=\hat{A}_{old}+R^{\star}(q,o)$ (9)
where $\hat{A}_{\text{old}}$ denotes the original advantage from outcome-based rewards. For PPO, we apply the exploration reward to the last advantage token. For GRPO, we apply it to each rollout advantage. This design preserves two key properties: (1) The outcome-based advantage maintains its original mean and variance statistics, ensuring stable policy updates; (2) Exploration rewards $R^{\star}(q,o)$ provide trajectory-level guidance without distorting value estimation. The decoupled formulation prevents exploration rewards from conflicting with outcome-based advantage normalization in GRPO or perturbing value function update in PPO, enabling harmonious integration of both reward components.
By applying trajectory-aware exploration rewards with dynamic reward scaling and advantage-preserving implementation, i-MENTOR efficiently enhances LLM reasoning by guiding the model toward diverse possible responses during training, while significantly improving overall reasoning ability. The complete algorithm is detailed in Appendix B.
3 Experiments
Here we evaluate i-MENTOR on three datasets to investigate the following research questions:
  • Q1: How does i-MENTOR enhance LLM reasoning performance in comparison with vanilla PPO and GRPO baselines?
  • Q2: What is the individual contribution of i-MENTOR’s core components to its overall effect?
In the following subsections, we first present the experimental setup, followed by a systematic analysis of results to address these questions. Notably, we also provide training statistics in Figure 3, a case study in Figure 2 and Appendix A, and comparisons with basic exploration techniques in Appendix D.
3.1 Experimental Setup
3.1.1 Dataset
We conduct experiments on three public datasets, i.e., GSM8K  [^2] https://huggingface.co/datasets/openai/gsm8k[9], Countdown-34  [^3] https://huggingface.co/datasets/Jiayi-Pan/Countdown-Tasks-3to4[11] and its harder version Countdown-4  [^4] https://huggingface.co/datasets/Jiayi-Pan/Countdown-Tasks-4[11] to validate i-MENTOR’s effectiveness against vanilla PPO and GRPO algorithms. For computational efficiency, we use a subset of the complete dataset of Countdown-34 and Countdown-4 for training. The detailed dataset statistic is shown in Appendix C.
3.1.2 Evaluation Protocol
Environment and LLM backbone: The experiment environment is built on the TinyZero  [^5] https://github.com/Jiayi-Pan/TinyZero[26] and verl  [^6] https://github.com/volcengine/verl framework with Qwen2.5-3B [44, 35] as base model. Similar to DAPO [46], we exclude the KL penalty from all RL algorithms after a detailed analysis in Appendix D.
Network structure of i-MENTOR: The predictor and target networks in i-MENTOR share the same architecture, which is dataset-agnostic and LLM-independent. For computational efficiency, we simply apply a lightweight predictor and target network structure for i-MENTOR (embedding size$=16$, three FFN layers with neurons $[16,8,1]$), ensuring negligible additional training time. To make a fairer comparison, we apply a set of fixed training hyperparameters to all three datasets.
Evaluation metric: For evaluation, we report the average accuracy on five experiments instead of evaluation scores to disable the potential gains brought by different format rewards during evaluation.
Additionally, we include implementation details in Appendix E to further elaborate on the training settings.
3.2 Main Result (Q1)
The main results are presented in Table 1. Additionally, Figure 4 highlights key training details via some of the training trajectories, showcasing the performance improvements achieved during training.
Table 1: Average accuracy on the three datasets. i-MENTOR-PPO denotes the implementation of i-MENTOR on PPO, while i-MENTOR-GRPO denotes the implementation of i-MENTOR on GRPO. “Improve” indicates relative improvement.
Dataset PPO i-MENTOR-PPO Improve GRPO i-MENTOR-GRPO Improve
GSM8K 0.8051 0.8169 1.47% 0.8082 0.8251 2.09%
Countdown-34 0.5526 0.5924 7.2% 0.6711 0.7132 6.27%
Countdown-4 0.3307 0.3812 15.27% 0.3872 0.4739 22.39%
(a) GSM8K with PPO.
(b) Countdown-34 with PPO.
(c) Countdown-4 with PPO.
(d) GSM8K with GRPO.
(e) Countdown-34 with GRPO.
(f) Countdown-4 with GRPO.
Figure 4: Evaluation accuracies in some of the training trajectories.
From Table 1 and Figure 4 we could conclude that:
  • GRPO outperforms PPO on all datasets, illustrating superior training effects brought by group-wise sampling and reward normalization. Both i-MENTOR-PPO and i-MENTOR-GRPO outperform standard RL approaches, demonstrating that i-MENTOR effectively guides and enhances the updates of LLM reasoning via exploration behavior during the RL training process. By encouraging the exploration of new responses rather than merely fitting to outcome-based reward through sampling behaviors, i-MENTOR enables LLMs to explore more diverse potential inference paths during training and avoids LLMs getting trapped in local optimal solutions.
  • A key challenge in improving LLM reasoning lies in handling difficult samples, which typically yield near-zero rewards that hinder parameter updates. Our experiments reveal an interesting pattern: i-MENTOR’s improvements are most significant on Countdown-4, followed by Countdown-34, while on GSM8K, where all RL methods converge within just 20 steps, the improvements are comparatively less pronounced. This progression aligns with the relative difficulty levels of these datasets, where Countdown-4 > Countdown-34 > GSM8K. The results suggest that i-MENTOR’s dense exploration rewards for each response sequence specifically help models overcome learning barriers in challenging samples.
  • i-MENTOR achieves greater performance gains with GRPO than with PPO. This advantage stems from GRPO’s rollout mechanism, which enables i-MENTOR to generate varied exploration rewards for multiple responses to the same question. Such a design promotes broader exploration of potential solution paths compared to single-response PPO updates.
3.3 Ablation Study (Q2)
The results from Table 2 highlight the progressive improvements contributed by each component of our proposed approach, i-MENTOR. Each enhancement effectively addresses core challenges in policy optimization for LLMs, as reflected by the corresponding increase in accuracy at every stage.
Table 2: Ablation study on Countdown-34.
Model Accuracy
GRPO 0.6711
$+$Trajectory-aware Exploration Rewards 0.6939
$+$Dynamic Reward Scaling: Exploration Reward Regularization 0.6988
$+$Dynamic Reward Scaling: Error-conditioned Reward Trigger 0.7007
$+$Dynamic Reward Scaling: Policy-preserving Exploration Attenuation 0.7065
$+$Advantage-Preserving Reward Implementation (i-MENTOR-GRPO) 0.7132
From the table, we could conclude that: (1) Incorporating trajectory-aware exploration rewards leads to a notable improvement in reasoning performance, increasing the accuracy from 0.6711 to 0.6939. This demonstrates that enabling the model to explore diverse potential responses during training promotes a richer learning process. By encouraging the discovery of alternative reasoning paths rather than converging prematurely on a limited set of high-reward responses, the model acquires more nuanced reasoning capabilities, as evidenced by the marked accuracy gain. (2) Adding the dynamic reward scaling mechanism further enhances model performance, with accuracy improving from 0.6939 to 0.7065. By applying exploration reward regularization, error-conditioned reward trigger, and policy-preserving exploration attenuation, i-MENTOR conducts efficient exploration with adaptive balance on exploration and exploitation during policy updates, ensuring stable and effective policy optimization. (3) Finally, the advantage-preserving reward implementation ensures seamless compatibility with RL algorithms, such as PPO and GRPO. By preserving the statistical integrity of outcome-based advantage distributions while incorporating exploratory signals, i-MENTOR achieves a robust balance between exploiting known strategies and exploring new areas, thus enhancing generalization capabilities. With all three components, i-MENTOR-GRPO demonstrates the highest improvement in accuracy, raising it to 0.7132.
Overall, the cumulative contributions of these techniques enable i-MENTOR to outperform the baseline GRPO model significantly, achieving a relative accuracy improvement of +6.27%. These results validate the effectiveness of our proposed components in addressing the unique challenges of training LLMs for complex reasoning tasks.
4 Related Works
4.1 Reinforcement Learning for LLMs
Reinforcement learning has evolved from foundational approaches like PPO [31] to advanced frameworks such as GRPO [33]. PPO achieves stability in policy updates through clipped objectives and trust region constraints, effectively aligning models for dialogue systems and code generation [38, 34]. Its architecture treats token positions in reasoning trajectories of LLM as distinct states for advantage estimation [16], though joint policy-value optimization introduces computational overhead [33].
GRPO [33] mitigates these limitations through trajectory-level sampling and group-wise advantage normalization. By generating multiple responses per input and standardizing rewards across trajectory groups, GRPO reduces policy update bias while maintaining efficiency. This sequence-level evaluation mechanism decouples advantage estimation from computational complexity, proving particularly effective in multi-step reasoning scenarios where traditional RL struggles with action space dimensionality. Recent implementations such as DAPO [46] and Dr. GRPO [21] further showcase GRPO’s scalability through detailed refinement of reasoning objectives, though exploration efficiency remains constrained by static and sparse reward structures [10, 12]. Differently, i-MENTOR enhances LLM reasoning through providing dense, stable, and efficient exploration rewards in RL optimization process. These rewards enable intrinsic motivation-guided exploration over diverse response trajectories during training, systematically improving reasoning capabilities.
4.2 Improving Reasoning Ability of LLMs
Recent advances in LLM reasoning focus on three key paradigms:
Pre-training augmentation improves foundational reasoning by exposing models to curated mathematical datasets and synthetic traces [45, 13, 33]. Models like Qwen [45] and Llama [13] show significant gains through domain-specific data scaling, though this requires heavy computational resources and careful data curation.
Prompt engineering [29, 7] enhances reasoning via structured inputs. Techniques like Chain-of-Thought [47, 8] decompose problems into steps, while self-consistency [39] improves outputs through majority voting. Extensions, such as multi-agent debates [18] and iterative refinement [23], further extend capabilities but are limited by manual design and high computational costs.
Algorithmic enhancement combines inference-time search with RL-based fine-tuning. Methods like Monte Carlo Tree Search [4, 42], beam search [43], and RL approaches (e.g., PPO [31] and GRPO [33]) optimize reasoning policies but often overexploit fixed high-reward paths, limiting exploration of novel solutions. Unlike these methods, our approach actively incentivizes exploration of under-optimized paths through exploration rewards, enabling deeper action-space traversal.
5 Conclusion
We present i-MENTOR, a method that addresses sparse rewards and enhances explorations in RL-based LLM reasoning optimization through three key advances. First, our trajectory-aware exploration reward mechanism uses lightweight neural networks to provide dense intrinsic rewards without adding computational overhead or episodic length bias, enabling efficient discovery of high-quality responses. Second, dynamic reward scaling automatically balances exploration and exploitation in large action spaces, ensuring sustained exploration while maintaining policy stability. Third, our advantage-preserving reward implementation decouples exploration incentives from outcome-based gradients, resolving conflicts between intrinsic and extrinsic rewards. Experiments show that i-MENTOR significantly improves reasoning performance with both vanilla PPO and GRPO, demonstrating its versatility and effectiveness. By stabilizing exploration-exploitation dynamics, i-MENTOR enables efficient training and diverse trajectory discovery in complex reasoning tasks.
Appendix A Detailed Case Study
We provide the full response texts from the case study (Figure 2) in Figure 5 for reference. Notably, the incorrect responses are longer than the correct ones, a common trend observed across all algorithms examined in this paper. This occurs because a well-trained LLM typically stops generating text once it arrives at the correct answer, whereas incorrect responses involve exploring more possible combinations to find a solution. However, as shown in Figure 3(a), the overall average response length of i-MENTOR-GRPO is longer than that of GRPO. This suggests that LLMs trained with i-MENTOR-GRPO develop more complex reasoning abilities, enabling them to approach problems from a broader range of perspectives.
Figure 5: Detailed case study text.
Appendix B i-MENTOR Algorithm
Through the methodological components described in Section 2.2, i-MENTOR implements a reinforcement learning framework with structured exploration guidance to enhance LLM reasoning capabilities. The complete optimization procedure, formalized in Algorithm 1, operates as follows: Given an input batch $B$ of question-response pairs $(q,o)$ for advantage estimation, we first compute the baseline outcome-based reward signals $R$ and initial advantage estimates $\hat{A}_{\text{old}}$ using conventional outcome-based reward functions and RL advantage estimation methods. Subsequently, each sample is processed through i-MENTOR’s dual networks $(f_{\theta_{P}},f_{\theta_{T}})$ to simultaneously update the policy network $f_{\theta_{P}}$ and obtain trajectory-aware exploration rewards $R^{\star}_{2}(q,o)$. These exploration rewards are then adaptively scaled through our dynamic reward scaling approaches, yielding $R^{\star}(q,o)$ that maintain training stability across different optimization phases. Finally, the refined advantages $\hat{A}_{\text{new}}$ are computed through our advantage-preserving implementation mechanism that injects $R^{\star}(q,o)$ into $\hat{A}_{\text{old}}$ without distorting the original policy gradient signals. The resulting $\hat{A}_{\text{new}}$ subsequently drives policy updates through standard RL optimization algorithms like PPO and GRPO, enabling effective LLM optimization while maintaining gradient stability.
Algorithm 1 Optimization algorithm of i-MENTOR
Input: A coming batch $B$ of $(q,o)$ samples for advantage estimation. A fixed randomly initialized network $f_{\theta_{T}}$, a predictor network $f_{\theta_{P}}$ with identical architecture as $f_{\theta_{T}}$.
Output: Advantage $\hat{A}_{new}$ for policy update with RL algorithms such as PPO and GRPO.
1:  Obtain outcome-based rewards $R$ through outcome-based reward function
2:  Obtain outcome-based advantage $\hat{A}_{old}$ via fixed reward functions
3:  Update $f_{\theta_{P}}$ according to loss function $\mathcal{L}$ in Equation (5)
4:  Obtain $R^{\star}_{2}(q,o)$ by predicting $r(q,o)$ and conduct min-max scaling in $B$ via Equation (6)
5:  Obtain $R^{\star}_{3}(q,o)$ from error-conditioned reward trigger via Equation (7)
6:  Obtain $R^{\star}(q,o)$ from policy-preserving exploration attenuation via Equation (8)
7:  Add the exploration reward $R^{\star}(q,o)$ to $\hat{A}_{old}$ for a new advantage $\hat{A}_{new}$ via Equation (9)
8:  return $\hat{A}_{new}$
Appendix C Dataset Statistics
This section briefly introduces the datasets used in this paper. Specifically, GSM8K 11footnotemark: 1 is a dataset of 8.5K high-quality linguistically diverse grade school math word problems. Countdown-34 22footnotemark: 2 and Countdown-4 33footnotemark: 3 are two mathematical datasets that perform combined operations based on several given numbers to obtain a given answer. Among them, the input sample of Countdown-34 contains 3 or 4 numbers, while the input sample of Countdown-4 only contains four numbers, making its average difficulty higher. For computational efficiency, we use a subset of the complete dataset of Countdown-34 and Countdown-4 for training. The detailed dataset statistic is shown in Table 3.
Table 3: Data Statistics.
Params GSM8K Countdown-34 Countdown-4
Training samples 7,473 32,768 32,768
Testing samples 1,319 1,024 1,024
Max prompt length 256 256 256
Max response length 1,024 1,024 1,024
Appendix D Comparison with Basic Exploration Techniques
Beyond i-MENTOR, researchers typically control LLM exploration through two basic techniques: (1) adjusting the temperature parameter $Temp$ to influence output diversity by reshaping token probabilities, and (2) modifying the KL penalty coefficient $\beta$ to regulate how strictly the policy adheres to its original behavior during RL updates. We evaluate these approaches using GRPO with varying KL and temperature coefficients, testing whether performance declines when deviating from their optimal values ($\beta=0.0$, $Temp=1.0$) used by default in this paper.
Our experiments in Table 4 reveal that increasing the KL coefficient (which tightens constraints on policy updates) and reducing the temperature (which introduces limited randomness in training) both degrade reasoning performance–the former limits exploratory updates by over-anchoring to the initial policy, while the latter disrupts sample diversity in training. This validates our baseline configuration with $\beta=0.0$ and $Temp=1.0$ for all RL methods in this paper. Moreover, by introducing exploration rewards that actively guide the model toward novel reasoning paths, i-MENTOR achieves superior performance. This demonstrates that i-MENTOR ’s trajectory-aware exploration rewards complement rather than conflict with basic exploration mechanisms, providing structured guidance for discovering novel responses while maintaining training stability.
Table 4: Comparison with Other Naive Exploration Methods. “$\beta$” and “Temperature” indicate the KL penalty and LLM temperature coefficients.
Model Countdown-4
GRPO ($\beta=0.0$, $Temp=1.0$) 0.3872
[…]
Appendix E Implementation Details
In this paper, we use 4 NVIDIA A800 GPUs for each training loop. Due to variations in dataset sizes and difficulty levels, we train for 250 steps on Countdown-3to4 and Countdown-4, and 20 steps on GSM8K, using a batch size of 512 to ensure convergence. Unless otherwise specified, all experiments are conducted with fixed hyperparameters for fair comparison. Based on preliminary grid search experiments, the exploration intensity parameter $\alpha$ is set to 0.5, and the exploration attenuation rate $\gamma$ is set to 40, ensuring an averaged optimal performance of i-MENTOR across all datasets. For GRPO, the rollout group size is set to 5. For outcome-based rewards, we adopt the same reward function as TinyZero 44footnotemark: 4, defined as:
$R(o,a)=\begin{cases}1.0,&a==o\\ 0.1,&\text{correct format reward}\\ 0.0,&\text{otherwise}\end{cases}$ (10)
where $a$ is the ground truth answer for response $o$. To ensure fair evaluation, we report the average accuracy over five experiments, rather than evaluation scores, to eliminate potential gains from format rewards during evaluation.
Appendix F Limitations
While i-MENTOR successfully encourages LLMs to explore novel responses, we believe that incorporating more diverse reward functions to evaluate the multi-faceted value of different responses and enabling multi-angle exploration could further enhance reasoning performance. As a future direction, we aim to better analyze these differences to guide LLMs in exploring large action spaces more effectively and developing more complex reasoning capabilities.
Appendix G Parameter Sensitivity Analysis
i-MENTOR involves two hyperparameters: $\alpha$ (Equation (6)), which scales the maximum exploration reward intensity, and $\gamma$ (Equation (8)), which regulates its decay rate. In this section, we visualize the parameter sensitivity experiments on Countdown-4 for i-MENTOR-GRPO in Figure 6 to analyze the influence of both on the performance of i-MENTOR.
Figure 6: Sensitivity of i-MENTOR to hyperparameters $\alpha$ and $\gamma$ on Countdown-4 for i-MENTOR-GRPO. Optimal performance occurs at $\alpha=0.5$, $\gamma=40$, adopted as default settings.
Our parameter sensitivity experiments and experiments on other datasets reveal three key insights: (1) Optimal $\alpha$-$\gamma$ combinations vary across datasets due to factors like task complexity. In this paper, $\alpha=0.5$ and $\gamma=40$ demonstrate robust performance as averaged defaults across the three datasets. (2) Extreme $\alpha$ values degrade performance–insufficient $\alpha$ limits exploration, while excessive $\alpha$ (especially with value $\gg 1$) risks overemphasizing exploration over correctness (e.g., exploration rewards surpassing the maximum value of outcome-based rewards in some samples). Notably, i-MENTOR with $\alpha=1.3$ and $\gamma=100$ still outperforms vanilla GRPO ($0.3872$), suggesting tolerance to moderate exploration emphasis. (3) Both slow ($\gamma\gg 40$) and rapid ($\gamma\ll 40$) decay rates harm performance: slow decay impedes training convergence, while rapid decay prematurely terminates exploration. This effect amplifies with larger $\alpha$ values.
These findings underscore i-MENTOR’s stability within practical parameter ranges while highlighting the necessity of balanced exploration-exploitation dynamics.